OpenStudy (anonymous):

Simply: 3[(5-3)2squared +(14 divide 2)squared]

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

is this the right interpretation? 3(5-3)^2 + (14/2)^2

OpenStudy (dumbcow):

\[3[(5-3)^{2} + (14/2)^{2}] = 3(2^{2} + 7^{2}) = 3(4+49) = 3(53) = 159\]